Circular No. 8483

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN CYGNUS S. Nakano, Sumoto, Japan, reports the discovery by Hideo Nishimura

(Kakegawa, Shizuoka-ken, Japan) of a possible nova (mag 9.7) on two films taken on Feb. 10.85 UT using a Pentax 6×7 camera (+ 200-mm f/4.0 lens + Kodak T-MAX 400 film; limiting mag 11). Nishimura provided the following semi-accurate position: α = 20^{h}09^{m}19^{s}, δ = +39^{o}48′49′′ (equinox 2000.0), adding that nothing was visible at this position on his patrol films back to 2001 Oct. 21 (including one taken on 2005 Feb. 6 with limiting mag 11). S. Wakuda (Yuto, Shizuoka-ken, Japan) provides position end figures 19^{s}.05, 52′′.9 (average of eight measurements) from a CCD image taken on Feb. 11.841 (the new object being at mag 9.3). K. Kadota (Ageo, Saitamaken, Japan) provides position end figures 19^{s}.09, 52′′.2 from an unfiltered image taken with a 0.25-m f/5.0 reflector on Feb. 11.761 (the object at mag 8.9). The nearest object in the USNO-A2.0 catalogue lies 14′′ to the southeast (and has red mag 16.4).

COMETS C/2004 Y12 AND C/2005 A2–A5 (SOHO) Additional Kreutz comets (cf. IAUC 8475; YT = Y.-s. Tsai):

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 Y12 Dec. 31.113 18^{h}57.^{m}3 −28^{o}09′ C3/2 TH 2005-C38
Comet  2005 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2005 A2 Jan. 3.404 19^{h}08.^{m}9 −25^{o}51′ C3 YT 2005-C38
C/2005 A3  6.488 19 23.4 −26 01 C3 HS 2005-C38
C/2005 A4  13.321 19 58.1 −25 03 C3 HS 2005-C38
C/2005 A5  15.405 20 01.0 −22 56 C3 RK 2005-C38

2005 AB V. Reddy, Department of Space Studies, University of North Dakota,

Grand Forks; R. Dyvig, Quinn, SD; and P. Pravec and P. Kušnirák, Ondřejov Observatory, report that photometric observations obtained during Feb. 1–8 show that the Amor-type minor planet 2005 AB (cf. MPEC 2005-A05, 2005-C48) is a binary system with orbital period 17.9 hr. The primary rotates with a period of 3.337 hr, and its lightcurve amplitude of 0.04 mag is indicative of a nearly spheroidal shape. Mutual events 0.06–0.11 mag deep indicate a secondary-to-primary mean-diameter ratio of 0.24 or greater. Further observations (especially radar and spectroscopic) are desirable.

2005 February 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green